# Trustworthy Transfer Learning: A Survey

### v1

@@ 3.2.3 Transferability on Time Series Data
Time series forecasting leverages historical observations to predict future events. As illustrated in Figure 6(a), there are two types of distribution shifts within time series forecasting. One is the sample-level temporal distribution shift (?, ?) of non-stationary time series where the data distribution of time series samples changes over time. The other one is the domain-level distribution shifts that occur between source and target time series domains (?). To address the first type of distribution shifts, AdaRNN (?) characterizes the distribution information by splitting the training sequences into diverse periods with the largest distribution gap, and then dynamically reduces the distribution discrepancy across these identified periods. RevIN (?) is a symmetrical normalization-and-denormalization method using instance normalization (?). It first normalizes the input sequences to mitigate distribution shifts among input sequences and then denormalizes the model outputs to restore the statistical information of input sequences. Follow-up approaches such as SAN (?) and Dish-TS (?) build upon RevIN to further address temporal distribution shifts between input and horizon sequences by adaptively learning normalization coefficients for fine-grained temporal slices. Additionally, SFA (?) incorporates test-time adaptation before forecasting to handle temporal distribution shifts. More recently, based on Koopman theory (?, ?), KNF (?) and Koopa (?) exploit linear Koopman operators to model the nonlinear dynamics of time series data on the measurement function space. Both methods design a global Koopman operator to learn time-invariant characteristics and a local Koopman operator to capture time-variant dynamics. To address the second type of distribution shifts, DAF (?) uses attention modules to learn complex temporal patterns within time series data and enforces the time-dependent query-key distribution alignment. Particularly, the queries and keys of attention modules are assumed to be domain-invariant, while the values capture domain-specific information for learning domain-dependent time series forecasters.
@@
Figure 6: Illustration of time series analysis under distribution shifts: (a) time series forecasting, and (b) time series classification
Definition 6 (Time Series Transferability for Classification (?))
@@
Given a source domain with labeled time series samples and a target domain with limited or no label information, time series transferability for classification aims to improve the prediction performance of the time series classification model in the target domain by leveraging knowledge from the source domain (shown in Figure 6(b)).
@@
Time series classification focuses on identifying time series data as a specific category. There are two main transfer learning frameworks for time series classification. The first framework involves pre-training a model on a source domain and then fine-tuning it on a target domain. For example, pre-training techniques for time series modeling have been developed using convolutional neural networks (?, ?), recurrent neural networks (?), ResNet (?, ?), and Transformers (?). The second framework is to learn domain-invariant time series representations using adversarial learning (?, ?, ?, ?, ?) or statistical divergence metrics (?, ?, ?, ?) using both source and target data. Specifically, VRADA (?) captures the domain-invariant temporal latent dependencies of multivariate time-series data using variational recurrent neural networks (?), followed by a gradient reversal layer (?). Similarly, CoDATS (?) is developed based on 1D convolutional neural networks and gradient reversal layer. CLUDA (?) and CALDA (?) further leverage contrastive learning losses to enhance the time series representations. In addition, AdvSKM (?) designs a hybrid spectral kernel network based on maximum mean discrepancy (?) to align source and target time series representations. Assuming that source and target time series domains share the same causal structure, SASA (?) uses long short-term memory (LSTM) networks to learn sparse associative structures from both domains, and then aligns them via maximum mean discrepancy (?). More recently, CauDiTS (?) furthermore employs an adaptive causal rationale disentanglement to learn domain-invariant causal rationales and domain-specific correlations from variable interrelationships. RAINCOAT (?) is proposed to address open-world adaptation scenarios, where source and target domains might have domain-specific private classes. It extracts both time features via 1-dimensional convolutional neural network and frequency features via discrete Fourier transform, and then aligns the time-frequency features across domains using Sinkhorn divergence (?).
4 Knowledge Trustworthiness
@@
In this section, we review the knowledge trustworthiness of transfer learning. Compared to standard trustworthy machine learning over a single domain, this survey discusses whether the source and target users can trust the transferred knowledge, whereas trustworthy machine learning investigates how a user can trust a model trained on private data. As illustrated in Figure 7, in the context of transfer learning, both the source domain owner and the target domain owner may have trustworthiness concerns about the transfer learning techniques. When considering the owners of the source domain as the “trustor”, do they trust that the transferred knowledge will not leak their data privacy (C1: Privacy)? Conversely, if the ”trustor” indicates the owners of the target domain, do they trust that the transferred knowledge is not poisoned (C2: Adversarial Robustness) or biased (C3: Fairness), and how well can the transferred knowledge be explained (C4: Transparency)?
Figure 7: Illustration of trustworthiness concerns in the knowledge transfer process
4.1 Privacy
@@
Privacy protection aims to prevent the unauthorized access or misuse of data that can directly or indirectly reveal sensitive private information, e.g., age, gender, login credential, fingerprint, medical records, etc. In recent years, privacy concerns in understanding the trustworthiness of artificial intelligence (AI) systems have been emphasized in released AI ethics guidelines (?, ?) and legal laws (e.g., General Data Protection Regulation (GDPR) (?) and California Consumer Privacy Act (CCPA) (?)). Maintaining privacy is critical in privacy-sensitive applications, such as patient clinical data analytics (?) and mobile keyboard prediction (?). Particularly, privacy protection in transfer learning frameworks focuses on preventing the leakage of private source data during the knowledge transfer process. This concern has inspired privacy-preserving transfer learning frameworks designed to transfer knowledge from a private source domain to a specific target domain while ensuring data privacy. One key principle of these frameworks is that all source data remains stored locally, with only the updated source models/hypotheses being shared securely.
4.1.1 Hypothesis Transfer
@@
Definition 7 (Hypothesis Transfer (?))
@@
$A^{\mathrm{htl}}:(\mathcal{X}\times\mathcal{Y})^{n_{T}}\times\mathcal{F}^{s}\to\mathcal{F}_{T}$ (16)
@@
As illustrated in Figure 8, there are three major learning scenarios for hypothesis transferability: (1) hypothesis transfer learning (?) with labeled target training data, (2) source-free adaptation (?) with unlabeled target training data, and (3) test-time adaptation (?) with only target testing data.
Figure 8: Illustration of source hypothesis transfer, including (a) hypothesis transfer learning with labeled target training data, (b) source-free adaptation with unlabeled target training data, and (c) test-time adaptation with only target testing data.
@@
    Hypothesis Transfer Learning: The goal of hypothesis transfer learning is to optimize the learning function on the target domain using the basis of hypotheses from the source domain (?). It assumes that both the source hypothesis and a few labeled target samples are accessible during the training of the target model. Earlier works (?, ?) utilized the source hypothesis as prior knowledge to guide the target learner in Bayesian learning frameworks. Following this, ? (?) theoretically analyze the generalization error (instantiated with the leave-one-out error) of regularized empirical risk minimization algorithms for hypothesis transfer learning. It is shown that the generalization error is positively correlated with a quantity $\mathbb{E}_{(x,y)\sim\mathbf{P}_{T}}[\ell(f_{S}(x),y)]$ measuring how the source hypothesis performs on the target domain, and hypothesis transfer learning enjoys faster convergence rates of generalization errors when a good source hypothesis is provided. Later, ? (?) extend generalization error bounds of regularized empirical risk minimization with (i) any non-negative smooth loss function, (ii) any strongly convex regularize, and (iii) a combination of multiple source hypotheses. They further highlight the impact of the quantity $\mathbb{E}_{(x,y)\sim\mathrm{P}_{T}}[\ell(f_{S}(x),y)]$ on the transfer performance and propose a principled approach to optimizing the combination of source hypotheses. Instead, ? (?) introduce a notion of transformation function to characterize the relatedness between the source and the target domains. Using this transformation function, they establish excess risk bounds for Kernel Smoothing and Kernel Ridge Regression. ? (?) further theoretically derive the optimal form of transformation functions under the squared loss scenario. More recently, ? (?) analyze hypothesis transfer learning through regularized empirical risk minimization under reproducing kernel Hilbert space (RKHS) with surrogate classification losses (e.g., exponential loss, logistical loss, softplus loss, mean squared error and squared hinge) in the context of binary classification. They establish the generalization error bounds and excess risk bounds based on hypothesis stability and pointwise hypothesis stability, highlighting the connections between surrogate classification losses and the quality of the source hypothesis. In addition, ? (?) and ? (?) study a more challenging few-shot hypothesis adaptation problem, where only a few labeled target samples (e.g., one sample per class) are available. Motivated by the learnability of semi-supervised learning, they propose generating highly-compatible unlabeled data to improve the training of the target learner.
@@
    It is noteworthy that gradient-based fine-tuning has become the predominant hypothesis transfer approach in the era of large foundation models (?, ?). Given a source hypothesis pre-trained on a source domain with adequate labeled samples, gradient-based fine-tuning aims to update the hypothesis through gradient descent optimization using a small amount of labeled target samples. The generalization performance of fine-tuning has been theoretically studied recently (?, ?). ? (?) show that the generalization error of fine-tuning under certain architectures (e.g., deep linear networks (?), shallow ReLU networks (?)) can be affected by the difference between optimal (normalized) source and target hypothesis, the covariance structure of the target data, and the depth of the network. Furthermore, recent works (?, ?, ?) show the generalization error of fine-tuning techniques in terms of the distance between fine-tuned and initialized model parameters.
@@
    Source-Free Adaptation: In contrast to hypothesis transfer learning, source-free adaptation (?, ?) enables hypothesis transfer from the source to the target domains using only unlabeled target data. To solve this problem, SHOT (?) is proposed to update the feature extractor of the pre-trained source model for the target domain while keeping the source classifier fixed. It maximizes the mutual information between intermediate feature representations and the output of the classifier, and also minimizes the prediction error using self-supervised pseudo labels. The follow-up works have developed source-free adaptation frameworks from various perspectives: clustering (?), pseudo-labeling (?, ?), data augmentation (?, ?), etc. As discussed in (?), most existing source-free adaptation approaches focus on understanding the discriminability-diversity trade-off: the former improves the discriminability of unlabeled target samples in the latent feature space while the latter ensures prediction diversity for all classes. In particular, ? (?) establish a theoretical connection between source-free adaptation and self-training (?) in terms of discriminability and diversity losses. This connection enables improved training of source-free adaptation by incorporating an auto-adjusting diversity constraint and teacher-student augmentation learning. In contrast, ? (?) and ? (?) study the discriminability-transferability trade-off in the context of source-free adaptation. Theoretically, ? (?) derive an information-theoretic generalization error bound for multi-source-free adaptation based on a bias-variance trade-off. Here, bias is triggered by the label and feature misalignments across domains, and variance is triggered by the number of pseudo-labeled target samples. ? (?) establish the connections between source-free adaptation and learning with noisy labels (?), given the findings that the pseudo-labels of target samples generated by the source model can be noisy due to domain shift. They theoretically justify the existence of the early-time training phenomenon (ETP) in source-free adaptation scenarios and propose using early learning regularization (?) to prevent the model from memorizing label noise during training. Empirically, in addition to standard vision tasks, ? (?) reevaluate existing source-free adaptation methods in a more challenging set of naturally occurring distribution shifts in bioacoustics. Their findings indicate that these existing methods often lack generalizability and perform worse than no adaptation in some cases. This highlights the necessity of evaluating source-free adaptation methods across a range of tasks, data modalities, and degrees of distribution shifts.
@@
    Test-Time Adaptation: Fully test-time adaptation (?) aims to adapt the source hypothesis to the target testing data, where data batches arrive sequentially and each batch can only be observed once. To this end, Test-Time Training (TTT) (?) and its modified version (TTT++) (?) incrementally update the feature extractor by minimizing the auxiliary task loss. Notably, this approach requires the optimize both this auxiliary task loss and the standard supervised loss. In contrast, without changing the training phase, Tent (?) is proposed to minimize the entropy of model predictions by only updating the normalization statistics and channel-wise affine transformations in an online manner. The follow-up works further improve this framework from two perspectives. One is to enhance the stability and robustness of test-time adaptation by minimizing the entropy of the average prediction across different augmentations (?, ?) or by minimizing sharpness-aware entropy (?, ?). The other addresses catastrophic forgetting by regularizing the updated parameters (?, ?) or adaptive resetting the model parameters (?). Furthermore, ? (?) analyze test-time adaptation through the lens of convex conjugate loss functions and propose a principled self-training approach based on conjugate pseudo labels for test-time adaptation. Later, ? (?) theoretically justify the advantages of conjugate labels over hard labels in test-time adaptation, by showing the performance gap between gradient descent with conjugate labels and gradient descent with hard labels in a binary classification problem. Empirically, ? (?) identify the commonly seen pitfalls when evaluating test-time adaptation algorithms, including sensitive hyperparameter selection, inconsistent source hypothesis, and insufficient consideration of various types of distribution shifts. ? (?) further demonstrate that the modules (e.g., batch normalization layers (?), feature extractor layers (?), classifier layers (?)) selected for test-time adaptation is strongly correlated with the types of distribution shifts.
4.1.2 Federated Transfer

### v2

@@ 3.2.3. Transferability on Time Series Data
Time series forecasting leverages historical observations to predict future events. As illustrated in Fig. 6(a), there are two types of distribution shifts within time series forecasting. One is the sample-level temporal distribution shift (RevIN; LiuWWL22) of non-stationary time series where the data distribution of time series samples changes over time. The other one is the domain-level distribution shifts that occur between source and target time series domains (DAF). To address the first type of distribution shifts, AdaRNN (AdaRNN) characterizes the distribution information by splitting the training sequences into diverse periods with the largest distribution gap, and then dynamically reduces the distribution discrepancy across these identified periods. RevIN (RevIN) is a symmetrical normalization-and-denormalization method using instance normalization (ulyanov2016instance). It first normalizes the input sequences to mitigate distribution shifts among input sequences and then denormalizes the model outputs to restore the statistical information of input sequences. Follow-up approaches such as SAN (SAN) and Dish-TS (Dish-TS) build upon RevIN to further address temporal distribution shifts between input and horizon sequences by adaptively learning normalization coefficients for fine-grained temporal slices. More recently, based on Koopman theory (koopman1931hamiltonian; brunton2022modern), KNF (KNF) and Koopa (Koopa) exploit linear Koopman operators to model the nonlinear dynamics of time series data on the measurement function space. Both methods design a global Koopman operator to learn time-invariant characteristics and a local Koopman operator to capture time-variant dynamics. To address the second type of distribution shifts, DAF (DAF) uses attention modules to learn complex temporal patterns within time series data and enforces the time-dependent query-key distribution alignment. Particularly, the queries and keys of attention modules are assumed to be domain-invariant, while the values capture domain-specific information for learning domain-dependent time series forecasters.
(a) Time series forecasting
(b) Time series classification
@@
Figure 6. Illustration of time series analysis under distribution shifts: (a) time series forecasting, and (b) time series classification
Definition 3.6 (Time Series Transferability for Classification (VRADA)).
@@
Given a source domain with labeled time series samples and a target domain with limited or no label information, time series transferability for classification aims to improve the prediction performance of the time series classification model in the target domain by leveraging knowledge from the source domain.
@@
Time series classification focuses on identifying time series data as a specific category (shown in Fig. 6(b)). There are two main transfer learning frameworks for time series classification. The first framework involves pre-training a model on a source domain and then fine-tuning it on a target domain. For example, pre-training techniques for time series modeling have been developed using convolutional neural networks (FawazFWIM18; ConvTimeNet), recurrent neural networks (TimeNet), ResNet (ZhangZTZ22; SimMTM), and Transformers (ZerveasJPBE21). The second framework is to learn domain-invariant time series representations using adversarial learning (VRADA; CoDATS; CLUDA; CALDA; CauDiTS) or statistical divergence metrics (SASA; OttRHBM22; AdvSKM; RAINCOAT) using both source and target data. Specifically, VRADA (VRADA) captures the domain-invariant temporal latent dependencies of multivariate time-series data using variational recurrent neural networks (VRNN), followed by a gradient reversal layer (DANN). Similarly, CoDATS (CoDATS) is developed based on 1D convolutional neural networks and a gradient reversal layer. CLUDA (CLUDA) and CALDA (CALDA) further leverage contrastive learning losses to enhance the time series representations. In addition, AdvSKM (AdvSKM) designs a hybrid spectral kernel network based on maximum mean discrepancy (MMD) to align source and target time series representations. Assuming that source and target time series domains share the same causal structure, SASA (SASA) uses long short-term memory (LSTM) networks to learn sparse associative structures from both domains, and then aligns them via maximum mean discrepancy (MMD). More recently, CauDiTS (CauDiTS) furthermore employs an adaptive causal rationale disentanglement to learn domain-invariant causal rationales and domain-specific correlations from variable interrelationships. RAINCOAT (RAINCOAT) is proposed to address open-world adaptation scenarios, where source and target domains might have domain-specific private classes. It extracts both time features via a 1-dimensional convolutional neural network and frequency features via the discrete Fourier transform, and then aligns the time-frequency features across domains using Sinkhorn divergence (Cuturi13).
4. Knowledge Trustworthiness
@@
In this section, we review the knowledge trustworthiness of transfer learning. Compared to standard trustworthy machine learning over a single domain, this survey discusses whether the source and target users can trust the transferred knowledge, whereas trustworthy machine learning investigates how a user can trust a model trained on private data. As illustrated in Fig. 7, in the context of transfer learning, both the source domain owner and the target domain owner may have trustworthiness concerns about the transfer learning techniques. When considering the owners of the source domain as the “trustor”, do they trust that the transferred knowledge will not leak their data privacy (C1: Privacy)? Conversely, if the ”trustor” indicates the owners of the target domain, do they trust that the transferred knowledge is not poisoned (C2: Adversarial Robustness) or biased (C3: Fairness), and how well can the transferred knowledge be explained (C4: Transparency)?
Figure 7. Illustration of trustworthiness concerns in the knowledge transfer process
4.1. Privacy
@@
Privacy protection aims to prevent the unauthorized access or misuse of data that can directly or indirectly reveal sensitive private information, e.g., age, gender, login credential, fingerprint, medical records, etc. In recent years, privacy concerns in understanding the trustworthiness of artificial intelligence (AI) systems have been emphasized in released AI ethics guidelines (jobin2019global; EuropeanGuidelines) and legal laws (e.g., General Data Protection Regulation (GDPR) (goodman2017european) and California Consumer Privacy Act (CCPA) (harding2019understanding)). Maintaining privacy is critical in privacy-sensitive applications, such as patient clinical data analytics (dayan2021federated) and mobile keyboard prediction (hard2018federated). Particularly, privacy protection in transfer learning frameworks focuses on preventing the leakage of private source data during the knowledge transfer process. This concern has inspired privacy-preserving transfer learning frameworks designed to transfer knowledge from a private source domain to a specific target domain while ensuring data privacy. One key principle of these frameworks is that all source data remains stored locally, with only the updated source models/hypotheses being shared securely.
4.1.1. Hypothesis Transfer
@@
Definition 4.1 (Hypothesis Transfer (KuzborskijO13)).
@@
(16) $A^{\mathrm{htl}}:(\mathcal{X}\times\mathcal{Y})^{n_{T}}\times\mathcal{F}^{s}\to\mathcal{F}_{T}$
@@
As illustrated in Fig. 8, there are three major learning scenarios for hypothesis transferability: (1) hypothesis transfer learning (KuzborskijO13) with labeled target training data, (2) source-free adaptation (SHOT) with unlabeled target training data, and (3) test-time adaptation (Tent) with only target testing data.
(a) Hypothesis transfer learning
(b) Source-free adaptation
(c) Fully test-time adaptation
Figure 8. Illustration of source hypothesis transfer, including (a) hypothesis transfer learning with labeled target training data, (b) source-free adaptation with unlabeled target training data, and (c) test-time adaptation with only target testing data.
@@
    Hypothesis Transfer Learning: The goal of hypothesis transfer learning is to optimize the learning function on the target domain using the basis of hypotheses from the source domain (KuzborskijO13). It assumes that both the source hypothesis and a few labeled target samples are accessible during the training of the target model. Earlier works (fei2006one; li2007bayesian) utilized the source hypothesis as prior knowledge to guide the target learner in Bayesian learning frameworks. (KuzborskijO13) theoretically analyze the generalization error (instantiated with the leave-one-out error) of regularized empirical risk minimization algorithms for hypothesis transfer learning. It is shown that the generalization error is positively correlated with a quantity $\mathbb{E}_{(x,y)\sim\mathbf{P}_{T}}[\ell(f_{S}(x),y)]$ measuring how the source hypothesis performs on the target domain, and hypothesis transfer learning enjoys faster convergence rates of generalization errors when a good source hypothesis is provided. Later, (kuzborskij2017fast) extend generalization error bounds of regularized empirical risk minimization with (i) any non-negative smooth loss function, (ii) any strongly convex regularizer, and (iii) a combination of multiple source hypotheses. They further highlight the impact of the quantity $\mathbb{E}_{(x,y)\sim\mathrm{P}_{T}}[\ell(f_{S}(x),y)]$ on the transfer performance and propose a principled approach to optimizing the combination of source hypotheses. Instead, (DuKSP17) introduce a notion of transformation function to characterize the relatedness between the source and the target domains. Using this transformation function, they establish excess risk bounds for Kernel Smoothing and Kernel Ridge Regression. (minami2023transfer) further theoretically derive the optimal form of transformation functions under the squared loss scenario. (AghbalouS23) analyze hypothesis transfer learning through regularized empirical risk minimization under reproducing kernel Hilbert space (RKHS) with surrogate classification losses (e.g., exponential loss, logistical loss, softplus loss, mean squared error, and squared hinge) in the context of binary classification. They establish the generalization error bounds and excess risk bounds based on hypothesis stability and pointwise hypothesis stability, highlighting the connections between surrogate classification losses and the quality of the source hypothesis. In addition, (TOHAN; DongLCLG0S023) study a more challenging few-shot hypothesis adaptation problem, where only a few labeled target samples (e.g., one sample per class) are available. Motivated by the learnability of semi-supervised learning, they propose generating highly-compatible unlabeled data to improve the training of the target learner.
@@
    It is noteworthy that gradient-based fine-tuning has become the predominant hypothesis transfer approach in the era of large foundation models (YosinskiCBL14; GoukHP21). Given a source hypothesis pre-trained on a source domain with adequate labeled samples, gradient-based fine-tuning aims to update the hypothesis through gradient descent optimization using a small amount of labeled target samples. The generalization performance of fine-tuning has been theoretically studied recently (ShachafBG21; JuLZ22). (ShachafBG21) show that the generalization error of fine-tuning under certain architectures (e.g., deep linear networks (JiT19), shallow ReLU networks (AroraDHLW19)) can be affected by the difference between optimal (normalized) source and target hypotheses, the covariance structure of the target data, and the depth of the network. Furthermore, recent works (GoukHP21; LiZ21; JuLZ22) show the generalization error of fine-tuning techniques in terms of the distance between fine-tuned and initialized model parameters.
@@
    Source-free Adaptation: In contrast to hypothesis transfer learning, source-free adaptation (KunduVVB20; SHOT) enables hypothesis transfer from the source to the target domains using only unlabeled target data. To solve this problem, SHOT (SHOT) is proposed to update the feature extractor of the pre-trained source model for the target domain while keeping the source classifier fixed. It maximizes the mutual information between intermediate feature representations and the output of the classifier, and also minimizes the prediction error using self-supervised pseudo labels. The follow-up works have developed source-free adaptation frameworks from various perspectives: clustering (yang2022attracting), pseudo-labeling (boudiaf2023search; lee2022confidence), data augmentation (KunduKBMKJR22; hwang2024sfda), etc. As discussed in (mitsuzumi2024understanding), most existing source-free adaptation approaches focus on understanding the discriminability-diversity trade-off: the former improves the discriminability of unlabeled target samples in the latent feature space while the latter ensures prediction diversity for all classes. In particular, (mitsuzumi2024understanding) establish a theoretical connection between source-free adaptation and self-training (wei2021theoretical) in terms of discriminability and diversity losses. This connection enables improved training of source-free adaptation by incorporating an auto-adjusting diversity constraint and teacher-student augmentation learning. In contrast, (KunduKBMKJR22; han2023discriminability) study the discriminability-transferability trade-off in the context of source-free adaptation. Theoretically, (shen2023balancing) derive an information-theoretic generalization error bound for multi-source-free adaptation based on a bias-variance trade-off. Here, bias is triggered by the label and feature misalignments across domains, and variance is triggered by the number of pseudo-labeled target samples. (yi2023when) establish the connections between source-free adaptation and learning with noisy labels, given the findings that the pseudo-labels of target samples generated by the source model can be noisy due to domain shift. They theoretically justify the existence of the early-time training phenomenon (ETP) in source-free adaptation scenarios and propose using early learning regularization (liu2020early) to prevent the model from memorizing label noise during training. Empirically, in addition to standard vision tasks, (boudiaf2023search) reevaluate existing source-free adaptation methods in a more challenging set of naturally occurring distribution shifts in bioacoustics. Their findings indicate that these existing methods often lack generalizability and perform worse than no adaptation in some cases. This highlights the necessity of evaluating source-free adaptation methods across a range of tasks, data modalities, and degrees of distribution shifts.
@@
    Test-time Adaptation: Fully test-time adaptation (Tent) aims to adapt the source hypothesis to the target testing data, where data batches arrive sequentially and each batch can only be observed once. To this end, Test-Time Training (TTT) (sun2020test) and its modified version (TTT++) (TTT++) incrementally update the feature extractor by minimizing the auxiliary task loss. Notably, this approach requires optimizing both this auxiliary task loss and the standard supervised loss. In contrast, without changing the training phase, Tent (Tent) is proposed to minimize the entropy of model predictions by only updating the normalization statistics and channel-wise affine transformations in an online manner. The follow-up works further improve this framework from two perspectives. One is to enhance the stability and robustness of test-time adaptation by minimizing the entropy of the average prediction across different augmentations (zhang2022memo) or by minimizing sharpness-aware entropy (niu2023towards; gong2023sotta). The other addresses catastrophic forgetting by regularizing the updated parameters (EATA; wang2022continual) or adaptive resetting the model parameters (niloy2024effective). Furthermore, (GoyalSRK22) analyze test-time adaptation through the lens of convex conjugate loss functions and proposes a principled self-training approach based on conjugate pseudo labels for test-time adaptation. Later, (wang2023towards) theoretically justify the advantages of conjugate labels over hard labels in test-time adaptation by showing the performance gap between gradient descent with conjugate labels and gradient descent with hard labels in a binary classification problem. Empirically, (zhao2023pitfalls) identify the commonly seen pitfalls when evaluating test-time adaptation algorithms, including sensitive hyperparameter selection, inconsistent source hypothesis, and insufficient consideration of various types of distribution shifts. (bao2023adaptive) further demonstrate that the modules (e.g., batch normalization layers (Tent), feature extractor layers (sun2020test), classifier layers (iwasawa2021test)) selected for test-time adaptation are strongly correlated with the types of distribution shifts.
4.1.2. Federated Transfer